OpenStudy (anonymous):

find the inverse of 2^x

OpenStudy (anonymous):

I figure i have to switch my x and y so \[x=2^{y}\]

OpenStudy (anonymous):

should I ln or log both sides or what?

pooja195 (pooja195):

Log(y) / Log(2)

zepdrix (zepdrix):

To get rid of the exponential base 2, you'll want to take the log base 2 of both sides :D it's a little awkward. But after you do that, you can do the log trick that pooja mentioned.

OpenStudy (anonymous):

so \[\log_{2}x=\log_{2}2^{y} \]

pooja195 (pooja195):

yep

OpenStudy (anonymous):

I dont get pooja's trick

OpenStudy (anonymous):

what should i do from there?

OpenStudy (anonymous):

well i know the right side would kinda cancel so i have \[\log_{2}x=y \]

zepdrix (zepdrix):

Changing the base of a log, here is example of the rule,\[\large \log_a (b)=\frac{\ln b}{\ln a}\]You can make the new base whatever you want, I did e in this case (the most natural of all logs :D)

OpenStudy (anonymous):

Im not seeing why you have to do that or what i'm really suposed to be doing. i"m kinda confused. sorry

zepdrix (zepdrix):

\[\large \log_2 x =\frac{\log_c x}{\log_c 2}\]It's a rule that lets us change our base to any new value \(c\) that we want. It's just a rule to remember :O I can't remember how it's derived actually...

OpenStudy (anonymous):

why do i need to change my base?

OpenStudy (anonymous):

i have \[\log_{2}x=y \]

zepdrix (zepdrix):

I suppose you don't! :) But the log base 2 is fairly insignificant. It's not a value that can be computed easily. Yah that's a fine answer ^^ prolly stop there. my bad

OpenStudy (anonymous):

oh so you are changing the base so the answer seems less trivial?

zepdrix (zepdrix):

Yah, which probably isn't necessary for your class XD hehe

OpenStudy (anonymous):

alright. awesome!